# Differentially Private Knowledge Distillation via Synthetic Text Generation

### v2

@@ 5.2 Main Results
For the DPKD baseline, we evenly allocated the privacy budget $\epsilon$ between the teacher and student, i.e., $T$ and $S$ use $\epsilon=1$ for DP-SGD. We observe that DPKD improves over the DP-SGD baseline by $2$, $0.6$ and $0.96$ PPL on the Yelp, Big Patent, DBpedia datasets, respectively. Hence, DPKD can give a modest performance boost to the student models compared to directly privately fine-tuning them for generative tasks, which was not the case for classification tasks shown in Mireshghallah et al. (2022).
@@
5010020040044444444464646464848484850505050Number of Synthetic Text Data (K)PPL
@@
One insightful observation is that for the Yelp and Big Patent datasets, DP Syn Data severely lags behind all baselines, only providing a modest improvement over the zero-shot baseline but not improving over the DP-SGD baseline. However, we do observe that for DBpedia, DP Syn Data can match the utility of DPKD. Additionally, we found that it is harder to optimize DP Syn Data due to the student easily overfitting the synthetic review data. This result highlights the fact that training on DP synthetically generated data alone is insufficient to competitively train a compressed DP student model for generative tasks. We surmise that a teacher model at the scale of a large foundation model is needed to generate high enough quality synthetic text to overcome this limitation Yu et al. (2023a).
@@ 7 Limitations
Another limitation is the same limitations from the synthetic text generation framework we used from Yue et al. (2022) carry over to our framework. In particular, small classes corresponding to control codes are disproportionately affected, where tight DP guarantees most negatively impact learning the distributions of small-size classes.
Finally, our work builds on recent advances in DP and LLM by initializing the models with pre-trained weights, and then privately fine-tuning them on a downstream dataset Li et al. (2021); Yu et al. (2021). However, recent work has shown that public LLMs can leak private information Carlini et al. (2021). A stronger and interesting future work would look to relax this assumption.
@@ Appendix A Additional Experimental Setup
Most of the hyperparameter values from table 1 were selected from Yue et al. (2022) and Mireshghallah et al. (2022), due to the large computational cost of hyperparameter searching of DP-SGD. Such values, like the clipping norm, are standard values to use. However, certain values were selected via hyperparameter sweeping, such as the learning rate. The computing infrastructure used to train the teacher and student models with DP-SGD is 8 40GB A100 GPUs, and the students with DP synthetic text data used 8 Quadro RTX 5000. The total training time for the teacher models on the Yelp dataset was 2 days and 17 hours, while on the Big Patent dataset, it took 10 hours. For the student models, it took about one days for the DP-SGD baseline, one and a half days for the DPKD baseline, and about five hours for DP Syn Data and ours. The reported main results are based on one run of training and then inference on the testing set. The version we used for the dp-transformers library Inan et al. (2022) is 1.0.1 which was installed by source.

### v3

@@ 5.2 Main Results
For the DPKD baseline, we evenly allocated the privacy budget $\epsilon$ between the teacher and student, i.e., $T$ and $S$ use $\epsilon=1$ for DP-SGD. We observe that DPKD improves over the DP-SGD baseline by $2$, $0.6$ and $0.96$ PPL on the Yelp, Big Patent, DBpedia datasets, respectively. Hence, DPKD can give a modest performance boost to the student models compared to directly privately fine-tuning them for generative tasks, which was not the case for classification tasks shown in mireshghallah2022differentially.
@@
501002004004444464648485050Number of Synthetic Text Data (K)PPL
@@
One insightful observation is that for the Yelp and Big Patent datasets, DP Syn Data severely lags behind all baselines, only providing a modest improvement over the zero-shot baseline but not improving over the DP-SGD baseline. However, we do observe that for DBpedia, DP Syn Data can match the utility of DPKD. Additionally, we found that it is harder to optimize DP Syn Data due to the student easily overfitting the synthetic review data. This result highlights the fact that training on DP synthetically generated data alone is insufficient to competitively train a compressed DP student model for generative tasks. We surmise that a teacher model at the scale of a large foundation model is needed to generate high enough quality synthetic text to overcome this limitation yu2023training.
@@ 7 Limitations
Another limitation is the same limitations from the synthetic text generation framework we used from yue2022synthetic carry over to our framework. In particular, small classes corresponding to control codes are disproportionately affected, where tight DP guarantees most negatively impact learning the distributions of small-size classes.
Finally, our work builds on recent advances in DP and LLM by initializing the models with pre-trained weights, and then privately fine-tuning them on a downstream dataset li2021large; yu2021differentially. However, recent work has shown that public LLMs can leak private information carlini2021extracting. A stronger and interesting future work would look to relax this assumption.
@@ Appendix A Additional Experimental Setup
Most of the hyperparameter values from table 1 were selected from yue2022synthetic and mireshghallah2022differentially, due to the large computational cost of hyperparameter searching of DP-SGD. Such values, like the clipping norm, are standard values to use. However, certain values were selected via hyperparameter sweeping, such as the learning rate. The computing infrastructure used to train the teacher and student models with DP-SGD is 8 40GB A100 GPUs, and the students with DP synthetic text data used 8 Quadro RTX 5000. The total training time for the teacher models on the Yelp dataset was 2 days and 17 hours, while on the Big Patent dataset, it took 10 hours. For the student models, it took about one days for the DP-SGD baseline, one and a half days for the DPKD baseline, and about five hours for DP Syn Data and ours. The reported main results are based on one run of training and then inference on the testing set. The version we used for the dp-transformers library inan2022dp-transformers is 1.0.1 which was installed by source.